OpenStudy (anonymous):

Simplify and leave your answer as a proper fraction or a mixed number 4 2/3+5 2/3

OpenStudy (anonymous):

\[4\frac{2}{3}+5\frac{2}{3} = \frac{8}{3}+\frac{10}{3}=\frac{8+10}{3}=\frac{18}{3}=6\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sorry my bad that is not right, i thought it was multiplication partial fraction answer is \[\frac{14}{3}+\frac{17}{3}=\frac{31}{3}=10\frac{1}{3}\]